# Modular degree and a conjecture of Watkins

### v2

@@ 2.3. Relations to congruence number $r_{E}$
Doi, Hida, Ribet, Mazur, and others investigated congruence primes [12, 34]; moreover, congruence primes were an essential component in Wiles’s [42] proof of Fermat’s last theorem.
Definition 1.
Let the newform linked to the elliptic curve $E$ be $f_{E}=\sum_{n>0}a_{n}q^{n}\in S_{2}\left(\Gamma_{0}(N)\right)$. The greatest integer such that there is a cuspform $g(z)=\sum_{n>0}b_{n}q^{n}$ with integer Fourier coefficients $b_{n}$, orthogonal to $f_{E}$ with respect to the Peterson inner product, and congruent to $f_{E}$ modulo $r_{E}$, is the congruence number $r_{E}$ of the elliptic curve $E$. (i.e., $a_{n}\equiv b_{n}\pmod{r_{E}}$ for all $n$).
Finding the link between congruence numbers and the modular degree is quite intriguing. Ribet showed that $m_{E}\mid r_{E}$. It was later extended by Amod, Ribet, and Stein [2] showing that if $\nu_{p}(N)\leq 1$, then $\nu_{p}(r_{E})=\nu_{p}(m_{E})$. Now one may naturally ask the following.
@@
Amod, Ribet, and Stein [2] establishes that when the conductor of an elliptic curve is squarefree, indicating that the curve is semistable, the values of ${r_{E}}$ and ${m_{E}}$ are equal. Abbes and Ullmo, in their work [1], have proven that if a prime number $p$ does not divide the conductor $N$, then $\nu_{p}(m_{E})=\nu_{p}(r_{E})$. Amod, Ribet, and Stein [2], in their conjecture, proposed that $\nu_{p}\left(\frac{r_{E}}{m_{E}}\right)\leq\frac{1}{2}\nu_{p}(N)$. In other words, they suggest that the order of $p$ dividing the ratio $\frac{r_{E}}{m_{E}}$ is at most half the order of $p$ dividing the conductor $N$. They also proved that $p$ is a prime such that $p\mid r_{E}$ but $p\nmid m_{E}$ then $\dim_{\mathbb{T}/\mathfrak{m}}J_{0}(N)[\mathfrak{m}]>2$, i.e., multiplicity one fails for $\mathfrak{m}$. Where $\mathbb{T}$ is the Hecke algebra acting on $E$, and $\mathfrak{m}$ be the annihilator of $E[p]$ in $\mathbb{T}$. The ideal $\mathfrak{m}$ is said to satisfies multiplicity one if $\dim_{\mathbb{T}/\mathfrak{m}}J_{0}(N)[\mathfrak{m}]=2$.
@@ 2.4. Congruence number of modular Abelian varieties
When $A$ is an elliptic curve, the modular exponent is equal to the modular degree, and the modular number is the square of the modular degree. For more details about the modular degree of abelian varieties and Watkins’s conjecture for abelian varieties, refer to the article by Dummigan and Krishnamoorthy[14]. Particularly [14, Corollary 9.5] discusses the $2$ power dividing the modular degree.
Definition 3.
@@
(4) $\frac{S_{2}(\Gamma_{0}(N),Z)}{S_{2}(\Gamma_{0}(N),Z)[I_{f}]\oplus S_{2}(\Gamma_{0}(N),Z)[I_{f}]^{\perp}}.$
@@
The natural question is to ask how the congruence number of Abelian varieties is related to the congruence number of elliptic curves when the Abelian variety is an elliptic curve. To answer this question, we state the following theorem, stated in [1] without proof. We provide a brief proof here.
Theorem 3.
Let $f_{E}=\sum_{n>0}a_{n}q^{n}\in S_{2}(\Gamma_{0}(N))$ be the newform linked to the elliptic curve $E$. Then the congruent number of the elliptic curve is consistent with Definition 3.
@@
By the above theorem, the congruence number of Abelian varieties is equal to the congruence number of elliptic curves when the Abelian variety is an elliptic curve. Amod, Ribet, and Stein [2] have also established analogous theorems to $m_{E}\mid r_{E}$ for modular Abelian varieties.
@@ 2.5. A lower bound on $2$-valuation
Let $E$ be a semistable curve with the conductor $N$. For each $d\mid\mid N$, let us consider the Atkin-Lehner involution $W_{d}$ on $X_{0}(N)$. The set $\left\{W_{d}|d\mid\mid N\right\}$ forms an abelian subgroup of automorphisms of rank equal to the number of prime dividing $N$. For every prime divisor $p\mid N$, we have $W_{p}(f_{E})=\pm f_{E}$. More precisely, there exists $w_{p}(f_{E})\in\{\pm 1\}$ such that $W_{p}(f_{E})=w_{p}(f_{E})f_{E}$. It turns out that $w_{p}(f_{E})=1$, if $E$ has non-split multiplicative reduction at $p$, and $-1$, if $E$ has split multiplicative reduction.
If $W_{p}(f_{E})=f_{E}$ Then the parameterization map $\phi:X_{0}(N)\to E$ factors through $X_{0}(N)/W_{p}$.
${X_{0}(N)}$${E}$${X_{0}(N)/W_{p}}$
The degree of the map $X_{0}(N)\to X_{0}(N)/W_{p}$ is $2$, contributing a factor of $2$ to the modular degree.
Let $\mu$ be the number of primes dividing $N$. Let us consider the homomorphism $(\mathbb{Z}/2\mathbb{Z})^{\mu}\to\{\pm 1\}$ defined by, $W_{d}\mapsto w_{d}$. Suppose $W^{\prime}$ is the kernel of this map. Since the field generated by the Fourier coefficients of the associated newform $f_{E}$ is $\mathbb{Q}$, it follows from Dummigan-Krishnamoorthy [14, Proposition 2.1] that there is a homomorphism $W^{\prime}\to E(\mathbb{Q})[2]$ with kernel $W^{\prime\prime}$, whose order divides the modular degree $m_{E}$. The order of $W^{\prime\prime}$ is at least greater than or equal to #$\frac{\#W^{\prime}}{E(\mathbb{Q})[2]}$. This gives,
$\nu_{2}(m_{E})$ $\geq\nu_{2}\left(\frac{\#W^{\prime}}{E(\mathbb{Q})[2]}\right)$
$=\nu_{2}\left({\#W^{\prime}}\right)-\nu_{2}\left({E(\mathbb{Q})[2]}\right)$
@@ 3. Divisibility and Watkins’s conjecture
In this section, we shall discuss the viability of Watkins’s conjecture. Recall from the introduction that Watkins’s conjecture predicts that $2^{\mathrm{rank}(E(\mathbb{(}Q)))}$ should divide $m_{E}$. We shall briefly discuss the approaches that have been taken to study this.
@@ 3.1. Heuristic proof of Watkins’s conjecture
Dummigan heuristically proved that certain classes of elliptic curves satisfy Watkins’s conjecture. Let $E$ be an elliptic curve with the square free even conductor $N$, with no rational point of order 2, and $E[2]$ is ramified at all primes $p\mid N$ and the action of complex conjugation on $E[2]$ is non-trivial then Dummigan [13] proved that $E$ satisfies Watkins’s conjecture assuming certain $R\cong T$. He uses Galois cohomology techniques and a squaring map to capture the weak Mordell-Weil group in a Selmer group. Later he established the isomorphism with the tangent space of universal deformation ring of residual representation $\overline{\rho}:{\mathrm{Gal}}(\overline{\mathbb{Q}}/{\mathbb{Q}})\to\mathrm{GL}_{2}(E[2])$ with certain deformation conditions. Local conditions in the $2$-Selmer group determine the deformation conditions. Later, using Wiles’s numerical criterion, he proved that $2^{\operatorname{rank}\left(E(\mathbb{Q})\right)}$ divides the modular degree of the elliptic curve.
@@
(5) $E_{A,B}:y^{2}=x^{3}+Ax+B,~{}A,B\in\mathbb{Z}~{}\mathrm{such~{}that}p^{6}~{}\mathrm{does~{}not~{}divide}~{}B~{}\mathrm{whenever}~{}p^{4}\mid A.$
@@
(6) $N(x)=\#\{A,B\in\mathbb{Z}:E_{A,B}\in\mathcal{E}(x),~{}\omega\left(\frac{4A^{3}+27B^{2}}{2^{\nu_{2}(4A^{3}+27B^{2})}3^{\nu_{3}(4A^{3}+27B^{2})}}\right)\leq 1\}=o(x).$
@@
(7) #$\#\{A,B\in\mathbb{Z}:|A|\leq x^{2},~{}|B|\leq x^{3}:p\mid 4B^{2}+27A^{3},~{}p>3\implies p>z\}=\frac{x^{5}}{\log z}+O(x^{2}z^{2}).$
Now, using (7), the quantity $N(x)$ at (6) can be estimated as follows
(8) $N(x)\leq\sum_{p<z}\#\{A,B\in\mathbb{Z}:E_{A,B}\in\mathcal{E}(x),~{}\frac{4A^{3}+27B^{2}}{2^{\nu_{2}(4A^{3}+27B^{2})}3^{\nu_{3}(4A^{3}+27B^{2})}}=\mathrm{a~{}power~{}of}~{}p\}+\frac{x^{5}}{\log z}+O(x^{2}z^{2}).$
@@
With this, we are now ready to prove Theorem 4
Proof of Theorem 4.
It follows from Lemma 1 that the conductor of almost the elliptic curves has at least two prime factors. Moreover, it follows from Grant’s result in [18] that almost all of them are $2$-torsion free over $\mathbb{Q}$. Now, the discussion in Section 2.5 implies that $\mathrm{Avg}_{2}(\mathrm{degree})\geq 1$. The proof from Bhargava, Shankar’s result in [6, Theorem 3] is complete. ∎
Remark 1.
Minimalist conjecture predicts that the average rank of elliptic curves should be 1/2; in particular, half of the curves have a rank 00, and half have rank $1$. Since Watkins’s conjecture is known to be true for all curves of rank $1$, we see that Watkins’s conjecture is true for almost all elliptic curves. This is a stronger version of Theorem 4.
@@ 3.2. Elliptic curves with odd modular degree
Calegari and Emerton undertook a significant endeavour to establish the validity of Watkins’s conjecture, focusing specifically on elliptic curves with odd modular degrees. Their approach involved developing a distinct characterization for elliptic curves falling within this category. Through an in-depth analysis of the Atkins-Lehner involution on $X_{0}(N)$, they effectively demonstrated that when an elliptic curve $E$ has an odd modular degree, the curve’s conductor $N$ is divisible by a maximum of two odd primes. Additionally, they established that the analytic rank of $E$ is necessarily even. They also show that if $E$ is an elliptic curve with an odd modular degree whose conductor $N$ has more than one prime factor, then $E$ possesses a rational point of order 2.
let $\mathbb{T}$ denote the Hecke algebra over $\mathbb{Z}$ acting on $S_{2}\left(\Gamma_{0}(N)\right)$, and let $\mathfrak{m}$ be a maximal ideal of $T$ such that $\mathbb{T}/\mathfrak{m}=\mathbb{F}_{2}$ Let $f_{E}$ be the eigenform of level $\Gamma_{0}(N)$ and weight $2$ corresponding to the elliptic curve $E$. According to the theorem in [2], the modular degree $m_{E}$ of $E$ is divisible by $2$ if and only if $f_{E}$ satisfies a congruence modulo $2$ with a cusp form of level $\Gamma_{0}(N)$. The set of cuspidal eigenforms congruent to $f_{E}$ is indexed by the set $\mathrm{Hom}(\mathbb{T}_{\mathfrak{m}}\times\mathbb{Q}_{2},\mathbb{Q}_{2})$. If $f_{E}$ does not satisfy any congruence with a cusp form, it means that there are no cuspidal eigenforms congruent to $f_{E}$. In other words, the set $\mathrm{Hom}(\mathbb{T}_{\mathfrak{m}}\times\mathbb{Q}_{2},\mathbb{Q}_{2})$ is trivial. Therefore, $f_{E}$ satisfies no congruence with a cusp form of level $\Gamma_{0}(N)$ if and only if $\mathbb{T}_{\mathfrak{m}}=\mathbb{Z}_{2}$. It is known that $\mathbb{T}_{\mathfrak{m}}=\mathbb{Z}_{2}$ equivalent to that there exists no nontrivial minimal deformation of ${\mathrm{GL}}_{2}(\mathbb{F}_{2}[x]/(x^{2}))$ of $\overline{\rho}$ , where $\overline{\rho}:{\mathrm{Gal}}(\overline{\mathbb{Q}}/{\mathbb{Q}})\to\mathrm{GL}_{2}(E[2])$ is the residual representation corresponding to $E[2]$.
@@
Subsequently, building upon the work of Calegari and Emerton [7] and Agashe et al. [2], Yazdani examined the elliptic curves with odd congruence numbers. They found that the conductors of all such elliptic curves, with a few exceptions, can be categorized as $p,pq,2p,4p$. Each of these categories was meticulously examined to demonstrate that all elliptic curves falling within these classifications possess a finite Mordell-Weil group, with the only potential exception being when the conductor is prime. In other words, Yazdani also proved that if an elliptic curve $E$ defined over $\mathbb{Q}$ has an odd modular degree, then its rank is zero unless it possesses a prime conductor and an even analytic rank. For a more comprehensive statement, please refer to [43, Theorem 3.8].
Furthermore, Kazalicki and Kohen [26] settled the case of elliptic curves with prime conductors by proving that those with an odd congruent number have a rank of zero. To establish their results, Kazalicki and Kohen utilized the Gross association, as described in Subsection 2.2, to establish a connection between the elliptic curve and an element of the Picard group. They then proved that if a rational elliptic curve $E$ has a prime conductor and a non-zero rank, the modular degree of $E$ is always even. Let us give a brief idea of the argument in the next section.

### v3

@@ 2.3. Relations to congruence number $r_{E}$
Doi, Hida, Ribet, Mazur, and others investigated congruence primes [DO76, R83]; moreover, congruence primes were an essential component in Wiles’s [Wiles95] proof of Fermat’s last theorem.
Definition 2.1.
Let the newform linked to the elliptic curve $E$ be $f_{E}=\sum_{n>0}a_{n}q^{n}\in S_{2}\left(\Gamma_{0}(N)\right)$. The greatest integer such that there is a cuspform $g(z)=\sum_{n>0}b_{n}q^{n}$ with integer Fourier coefficients $b_{n}$, orthogonal to $f_{E}$ with respect to the Petersson inner product, and congruent to $f_{E}$ modulo $r_{E}$, is the congruence number $r_{E}$ of the elliptic curve $E$. (i.e., $a_{n}\equiv b_{n}\pmod{r_{E}}$ for all $n$).
Finding the link between congruence numbers and the modular degree is quite intriguing. Ribet showed that $m_{E}\mid r_{E}$. It was later extended by Amod, Ribet, and Stein [ARS] showing that if $\nu_{p}(N)\leq 1$, then $\nu_{p}(r_{E})=\nu_{p}(m_{E})$. Now one may naturally ask the following.
@@
Amod, Ribet, and Stein [ARS] establishes that when the conductor of an elliptic curve is squarefree, indicating that the curve is semistable, the values of $r_{E}$} and $m_{E}$} are equal. Abbes and Ullmo, in their work [AU96], have proven that if a prime number $p$ does not divide the conductor $N$, then $\nu_{p}(m_{E})=\nu_{p}(r_{E})$. Amod, Ribet, and Stein [ARS], in their conjecture, proposed that $\nu_{p}\left(\frac{r_{E}}{m_{E}}\right)\leq\frac{1}{2}\nu_{p}(N)$. In other words, they suggest that the order of $p$ dividing the ratio $\frac{r_{E}}{m_{E}}$ is at most half the order of $p$ dividing the conductor $N$. They also proved that $p$ is a prime such that $p\mid r_{E}$ but $p\nmid m_{E}$ then $\dim_{\mathbb{T}/\mathfrak{m}}J_{0}(N)[\mathfrak{m}]>2$, i.e., multiplicity one fails for $\mathfrak{m}$. Where $\mathbb{T}$ is the Hecke algebra acting on $E$, and $\mathfrak{m}$ be the annihilator of $E[p]$ in $\mathbb{T}$. The ideal $\mathfrak{m}$ is said to satisfies multiplicity one if $\dim_{\mathbb{T}/\mathfrak{m}}J_{0}(N)[\mathfrak{m}]=2$.
@@ 2.4. Congruence number of modular Abelian varieties
When $A$ is an elliptic curve, the modular exponent is equal to the modular degree, and the modular number is the square of the modular degree. For more details about the modular degree of abelian varieties and Watkins’s conjecture for abelian varieties, refer to the article by Dummigan and Krishnamoorthy[DK13]. Particularly [DK13, Corollary 9.5] discusses the $2$ power dividing the modular degree.
Definition 2.3.
@@
(2.4) $\frac{S_{2}(\Gamma_{0}(N),Z)}{S_{2}(\Gamma_{0}(N),Z)[I_{f}]\oplus S_{2}(\Gamma_{0}(N),Z)[I_{f}]^{\perp}}.$
@@
The natural question is to ask how the congruence number of Abelian varieties is related to the congruence number of elliptic curves when the Abelian variety is an elliptic curve. To answer this question, we state the following theorem, stated in [AU96] without proof. We provide a brief proof here.
Theorem 2.4.
Let $f_{E}=\sum_{n>0}a_{n}q^{n}\in S_{2}(\Gamma_{0}(N))$ be the newform linked to the elliptic curve $E$. Then the congruent number of the elliptic curve is consistent with Definition 2.3.
@@
By the above theorem, the congruence number of Abelian varieties is equal to the congruence number of elliptic curves when the Abelian variety is an elliptic curve. Amod, Ribet, and Stein [ARS] have also established analogous theorems to $m_{E}\mid r_{E}$ for modular Abelian varieties.
@@ 2.5. A lower bound on $2$-valuation
Let $E$ be a semistable curve with the conductor $N$. For each $d\mid\mid N$, let us consider the Atkin-Lehner involution $W_{d}$ on $X_{0}(N)$. The set $\left\{W_{d}|d\mid\mid N\right\}$ forms an abelian subgroup of automorphisms of rank equal to the number of prime dividing $N$. For every prime divisor $p\mid N$, we have $W_{p}(f_{E})=\pm f_{E}$. More precisely, there exists $w_{p}(f_{E})\in\{\pm 1\}$ such that $W_{p}(f_{E})=w_{p}(f_{E})f_{E}$. It turns out that $w_{p}(f_{E})=1$, if $E$ has non-split multiplicative reduction at $p$, and 1-1, if $E$ has split multiplicative reduction. If $W_{p}(f_{E})=f_{E}$ Then the parameterization map $\phi:X_{0}(N)\to E$ factors through $X_{0}(N)/W_{p}$.
[…]
The degree of the map $X_{0}(N)\to X_{0}(N)/W_{p}$ is $2$, contributing a factor of $2$ to the modular degree. Let $\mu$ be the number of primes dividing $N$. Let us consider the homomorphism $(\mathbb{Z}/2\mathbb{Z})^{\mu}\to\{\pm 1\}$ defined by, $W_{d}\mapsto w_{d}$. Suppose $W^{\prime}$ is the kernel of this map. Since the field generated by the Fourier coefficients of the associated newform $f_{E}$ is $\mathbb{Q}$, it follows from Dummigan-Krishnamoorthy [DK13, Proposition 2.1] that there is a homomorphism $W^{\prime}\to E(\mathbb{Q})[2]$ with kernel $W^{\prime\prime}$, whose order divides the modular degree $m_{E}$. The order of $W^{\prime\prime}$ is at least greater than or equal to $\frac{\#W^{\prime}}{E(\mathbb{Q})[2]}$. This gives,
$\nu_{2}(m_{E})$ $\geq\nu_{2}\left(\frac{\#W^{\prime}}{\#E(\mathbb{Q})[2]}\right)$
$=\nu_{2}\left({\#W^{\prime}}\right)-\nu_{2}\left(\#E(\mathbb{Q})[2]\right)$
@@ 3. Divisibility and Watkins’s conjecture
In this section, we shall discuss the viability of Watkins’s conjecture. Recall from the introduction that Watkins’s conjecture predicts that $2^{\mathrm{rank}(E(\mathbb{Q}))}$ should divide $m_{E}$. We shall briefly discuss the approaches that have been taken to study this.
@@ 3.1. Heuristic proof of Watkins’s conjecture
Dummigan heuristically proved that certain classes of elliptic curves satisfy Watkins’s conjecture. Let $E$ be an elliptic curve with the squarefree even conductor $N$, with no rational point of order 2, and $E[2]$ is ramified at all primes $p\mid N$ and the action of complex conjugation on $E[2]$ is non-trivial, then Dummigan [neil06] proved that $E$ satisfies Watkins’s conjecture assuming certain $R\cong T$. He used Galois cohomology techniques and a squaring map to capture the weak Mordell-Weil group in a Selmer group. Later he established the isomorphism with the tangent space of the universal deformation ring of residual representation $\overline{\rho}:{\mathrm{Gal}}(\overline{\mathbb{Q}}/{\mathbb{Q}})\to\mathrm{GL}_{2}(E[2])$ with certain deformation conditions. Local conditions in the $2$-Selmer group determine the deformation conditions. Later, using Wiles’s numerical criterion, he proved that $2^{\operatorname{rank}\left(E(\mathbb{Q})\right)}$ divides the modular degree of the elliptic curve.
@@
(3.1) $E_{A,B}:y^{2}=x^{3}+Ax+B,~A,B\in\mathbb{Z}~\mathrm{such~that}p^{6}~\mathrm{does~not~divide}~B~\mathrm{whenever}~p^{4}\mid A.$
@@
(3.2) $N(x)=\#\{A,B\in\mathcal{E}(x):\omega\left(\frac{4A^{3}+27B^{2}}{2^{\nu_{2}(4A^{3}+27B^{2})}3^{\nu_{3}(4A^{3}+27B^{2})}}\right)\leq 1\}=o(x^{5}).$
@@
(3.3) $\{A,B\in\mathbb{Z}:|A|\leq x^{2},~|B|\leq x^{3}:p\mid 4B^{2}+27A^{3},~p>3\implies p>z\}=\frac{x^{5}}{\log z}+O(x^{2}z^{2}).$
Now, using (3.3), the quantity $N(x)$ at (3.2) can be estimated as follows
(3.4) $N(x)\leq\sum_{p<z}\#\{A,B\in\mathbb{Z}:E_{A,B}\in\mathcal{E}(x),~\frac{4A^{3}+27B^{2}}{2^{\nu_{2}(4A^{3}+27B^{2})}3^{\nu_{3}(4A^{3}+27B^{2})}}=\mathrm{a~power~of}~p\}+\frac{x^{5}}{\log z}+O(x^{2}z^{2}).$
@@
With this, we are now ready to prove Theorem 3.1
Proof of Theorem 3.1.
It follows from Lemma 3.2 that the conductor of almost the elliptic curves has at least two prime factors. Moreover, it follows from Grant’s result in [grant00] that almost all of them are $2$-torsion free over $\mathbb{Q}$. Now, the discussion in Section 2.5 implies that $\mathrm{Avg}_{2}(\mathrm{degree})\geq 1$. The proof from Bhargava, Shankar’s result in [BS13, Theorem 3] is complete. ∎
Remark 3.3.
Minimalist conjecture predicts that the average rank of elliptic curves should be 1/2; in particular, half of the curves have a rank 0, and half have rank $1$. Since Watkins’s conjecture is known to be true for all curves of rank $1$, we see that Watkins’s conjecture is true for almost all elliptic curves. This is a stronger version of Theorem 3.1.
@@ 3.2. Elliptic curves with odd modular degree
Calegari and Emerton undertook a significant endeavour to establish the validity of Watkins’s conjecture, focusing specifically on elliptic curves with odd modular degrees. Their approach involved developing a distinct characterization for elliptic curves falling within this category. Through an in-depth analysis of the Atkin–Lehner involution on $X_{0}(N)$, they effectively demonstrated that when an elliptic curve $E$ has an odd modular degree, the curve’s conductor $N$ is divisible by a maximum of two odd primes. Additionally, they established that the analytic rank of $E$ is necessarily even. They also show that if $E$ is an elliptic curve with an odd modular degree whose conductor $N$ has more than one prime factor, then $E$ possesses a rational point of order 2.
Let $\mathbb{T}$ denote the Hecke algebra over $\mathbb{Z}$ acting on $S_{2}\left(\Gamma_{0}(N)\right)$, and let $\mathfrak{m}$ be a maximal ideal of $\mathbb{T}$ such that $\mathbb{T}/\mathfrak{m}=\mathbb{F}_{2}$. Let $f_{E}$ be the eigenform of level $\Gamma_{0}(N)$ and weight $2$ corresponding to the elliptic curve $E$. According to the theorem in [ARS], the modular degree $m_{E}$ of $E$ is divisible by $2$ if and only if $f_{E}$ satisfies a congruence modulo $2$ with a cusp form of level $\Gamma_{0}(N)$. The set of cuspidal eigenforms congruent to $f_{E}$ is indexed by the set $\mathrm{Hom}(\mathbb{T}_{\mathfrak{m}}\times\mathbb{Q}_{2},\mathbb{Q}_{2})$. If $f_{E}$ does not satisfy any congruence with a cusp form, it means that there are no cuspidal eigenforms congruent to $f_{E}$. In other words, the set $\mathrm{Hom}(\mathbb{T}_{\mathfrak{m}}\times\mathbb{Q}_{2},\mathbb{Q}_{2})$ is trivial. Therefore, $f_{E}$ satisfies no congruence with a cusp form of level $\Gamma_{0}(N)$ if and only if $\mathbb{T}_{\mathfrak{m}}=\mathbb{Z}_{2}$. It is known that $\mathbb{T}_{\mathfrak{m}}=\mathbb{Z}_{2}$ equivalent to that there exists no nontrivial minimal deformation of ${\mathrm{GL}}_{2}(\mathbb{F}_{2}[x]/(x^{2}))$ of $\overline{\rho}$ , where $\overline{\rho}:{\mathrm{Gal}}(\overline{\mathbb{Q}}/{\mathbb{Q}})\to\mathrm{GL}_{2}(E[2])$ is the residual representation corresponding to $E[2]$.
@@
Subsequently, building upon the work of Calegari and Emerton [CE09] and Agashe et al. [ARS], Yazdani examined the elliptic curves with odd congruence numbers. They found that the conductors of all such elliptic curves, with a few exceptions, can be categorized as $p,pq,2p,4p$. Each of these categories was meticulously examined to demonstrate that all elliptic curves falling within these classifications possess a finite Mordell-Weil group, with the only potential exception being when the conductor is prime. In other words, Yazdani also proved that if an elliptic curve $E$ defined over $\mathbb{Q}$ has an odd modular degree, then its rank is zero unless it possesses a prime conductor and an even analytic rank. For a more comprehensive statement, please refer to [yazdani11, Theorem 3.8].
Furthermore, Kazalicki and Kohen [KK18watkinscorrigendum] settled the case of elliptic curves with prime conductors by proving that those with an odd congruent number have a rank of zero. To establish their results, Kazalicki and Kohen utilized the Gross association, as described in Subsection 2.2, to establish a connection between the elliptic curve and an element of the Picard group. They then proved that if a rational elliptic curve $E$ has a prime conductor and a non-zero rank, the modular degree of $E$ is always even. Let us give a brief idea of the argument in the next section.